OpenStudy (anonymous):

If possible, solve by factoring the polynomial equation. Show all work. I will award a medal and become a fan! x^3 + x^2 - 25x -25 = 0

OpenStudy (anonymous):

Help.

OpenStudy (chosenmatt):

Simplifying x3 + x2 + -25x + -25 = 0 Reorder the terms: -25 + -25x + x2 + x3 = 0 Solving -25 + -25x + x2 + x3 = 0 now go from there

OpenStudy (anonymous):

I don't think there are any GCF's here.